OpenStudy (anonymous):

3(x+3)=24 solve the equation

OpenStudy (anonymous):

divide by3 u get x+3=8--->x=5

OpenStudy (anonymous):

3(x+3)=24 multiply 3 to (x+3)= 3x+9=24 transpose 9 to the other side 3x=24-9 3x=15 divide both side on 3 x=5